OpenStudy (alyssakin):

Two kids are roller skating. Amy, with a mass of 55 kg, is traveling forward at 3 m/s. Jenny, who has a mass of 40 kg, is traveling in the opposite direction at 5 m/s. They crash into each other and hold onto each other so that they move as one mass. How fast are they traveling? I'm so confused please help.

OpenStudy (onepieceftw):

velocity = mass * acceleration

OpenStudy (welshfella):

this is an exercise conservation of momentum The total momentum before the crash = total after the crash. Momentum = mass * velocity

OpenStudy (alyssakin):

Thanks.

OpenStudy (welshfella):

note that momentum is a vector so that if you take Jenny's velocity to be 5 m/s then Amy's velocity will be -3 m/s

OpenStudy (welshfella):

do you need any more help?

OpenStudy (alyssakin):

Hold on, I'm gonna try to answer it and then I'll ask if I did it right.

OpenStudy (welshfella):

fine

OpenStudy (alyssakin):

Okay. So Amy 55 * 3 = 220 N Jenny 40 * 5 = 200 S Do I subtract those two to get the answer?

OpenStudy (welshfella):

55 * 3 = 165

OpenStudy (alyssakin):

Oops. Thanks for catching that.

OpenStudy (welshfella):

You add these 2 values 220 + (-165) this is the totalmomentum before the collision After the collision the mass of the 2 girls = 85 kg so that the momentum is 95v where v is their velocity So equate these 2 and solve for v

OpenStudy (welshfella):

* Sorry that's 95 kg not 85

OpenStudy (alyssakin):

Okay. Thanks a ton!!!

OpenStudy (welshfella):

yw